# Are foundation models for computer vision good conformal predictors?

### v2

@@ 4.2.1 Performance in the General Setting
First, we study the performance of 11 vision foundation models paired with CP methods under the standard setting, and on the three datasets, which presents ideal conditions: a sufficiently large calibration set, and absence of distributional drifts between calibration and test sets. We aim to determine whether we can prescribe a winner solution in this scenario and which factors can help identifying it.
@@
Fig. 1 depicts the relationship between the linear probing performance of each model and the conformal set size (top), and coverage gap (bottom). The initial observation highlights a clear trend: higher-performance models tend to produce smaller prediction sets, regardless of the conformal method used. Nevertheless, while set efficiency (i.e. size) is typically considered as a sufficient condition in most prior literature in conformal prediction [1], our analysis revealed that the relation with the accuracy does not consistently hold when examining other metrics. In particular, Fig. 1 (bottom) exposes that different CP methods yield mixed results for the coverage gap, which do not correlate directly with accuracy across all CP approaches. While APS appears to be almost unaffected by the model performance, RAPS clearly benefits from a strong performance, and LAC is negatively affected by more accurate models as the dataset becomes more complex.
@@
Regarding comparison between the CP methods, if we consider the set size, APS is clearly outperformed by the other approaches, particularly on ImageNet, whereas LAC provides the smallest prediction sizes, closely followed by RAPS. Indeed, RAPS is specifically designed to reduce the conformal set size of APS. However, the coverage gap results indicate that this comes at the cost of increasing the range of the class-conditional coverage. Below, we analyze the underlying causes that may explain this behaviour.
@@
RAPS class-conditional coverage, and therefore coverage gap, are more sensitive to the model’s accuracy. Let us assume we have two models, $M_{1}$ and $M_{2}$, in a multi-class classification problem, whose accuracies are $Acc_{1}$ and $Acc_{2}$, respectively. For each class $y\in\mathcal{Y}$, we refer to $C_{M_{i}}(y)=\mathbb{P}(Y=y|Y\in\mathcal{S}_{M_{i}}(X))$ as the class-conditional coverage for $y$ under model $M_{i}$, which measures the probability that predictions include the true label when the true label is $y$. Furthermore, let $\delta_{M_{1}}=\min_{y\in\mathcal{Y}}C_{M_{1}}(y)$ and $\delta_{M_{2}}=\min_{y\in\mathcal{Y}}C_{M_{2}}(y)$ denote the minimum class-conditional coverage achieved by each model. Under this scenario, we argue that due to the penalty in RAPS, if $Acc_{1}<Acc_{2}$, then $\delta_{M_{1}}<\delta_{M_{2}}$ when using RAPS as a conformal prediction method. In particular, $M_{1}$ (with lower accuracy) needs to expand its prediction sets for certain classes to meet the marginal coverage target 1-$\alpha$. However, the penalty term encourages small prediction sets, limiting an excessive number of classes. Thus, for some difficult classes, model $M_{1}$ may still potentially fail to meet the target coverage, as the enforced penalty discourages overly large sets. This ultimately results in lower coverage rates for those specific classes than $M_{2}$. In contrast, since APS does not include any regularization term that encourages small set sizes, it will compensate for lower performing models by increasing its set sizes, ultimately attaining higher class-conditional coverages.
@@
Figure 2: Comparison (APS vs RAPS) of the class-conditional coverage and set size for the class for which RAPS has the worst class-conditional coverage. Experiments performed on CIFAR-100. Models sorted (in ascending order) by their LP performance ($\min=0.65$ and $\max=0.92$), indicated by the size of the circles.
@@
To confirm this hypothesis, we perform the following experiment. First, we identify the class with the lowest minimum class-conditional coverage obtained by RAPS for each model $m$, which we refer to as $y^{R}_{m}$, and find its corresponding set size (both represented blue circles in Fig. 2). Then, we identify the class-conditional coverage and set size of the class $y^{R}_{m}$ provided by APS across all models, whose values are shown as pink circles in both plots of Fig. 2. Last, we also include the minimum class-conditional coverage obtained by APS across each method, depicted with a cross. Note that the minimum class-conditional coverage from APS does not necessarily correspond to the coverage of class $y^{R}_{m}$. Upon close examination of these results, we observe that, indeed, for models presenting lower accuracy (those with smaller circles), the gap between the class-conditional coverage for the worst class in RAPS and the same class in APS is consistently larger than for more accurate models (larger circles). This empirical evidence shows that the minimum class-conditional coverage is substantially reduced because RAPS is constrained from expanding its predictions set size. This effect is more pronounced in less accurate models, where the true class may rank far away from the maximum allowable set size in the softmax predictions. In contrast, APS tolerates worse models by increasing the set size, which ultimately degrades the set efficiency but yields better class-conditional coverages.
@@
Set size ($\downarrow$) MCCC ($\uparrow$)
Acc ($\uparrow$) Thr APS RAPS Thr APS RAPS
$\text{ViT}_{\text{CLIP}}$ 72.01 3.03 9.50 3.73 0.434 0.556 0.418
$\text{ViT}_{\text{DINO-S}}$ 74.92 3.27 10.02 4.23 0.433 0.477 0.412
$\text{ViT}_{\text{MetaCLIP}}$ 75.80 2.39 9.43 2.84 0.479 0.535 0.467
$\text{ViT}_{\text{ImageNet}}$ 76.08 2.36 38.75 4.46 0.416 0.495 0.405
@@
Following this analysis, we are also interested in determining whether a network pre-trained following a more traditional approach (i.e., standard supervised fine-tuning) offers similar conformal capabilities to self-supervised and contrastive ones. In particular, we select a ViT-B pre-trained on ImageNet, which is the same architecture as the visual encoder of the different foundation models. The results from this analysis (Table 1) reveal that, despite obtaining lower classification accuracy when using LP on the different foundation models, CP methods typically yield better performance than in $\text{ViT}_{\text{ImageNet}}$. These differences are significant under the APS approach, where the set size is significantly degraded on $\text{ViT}_{\text{ImageNet}}$. Moreover, the class-conditional coverage is also substantially affected, with nearly 6% decrease compared to the best model.
@@
Figure 4: Evaluation under domain-shift. Set size($\downarrow$), coverage($\uparrow$), and MCCC($\uparrow$) across three CP methods and three foundation models. ImageNet versions are sorted based on OOD performance in [53].
To further delve into these differences, we compute, for each test sample, the difference between the conformal set size for APS when applied to $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{CLIP}}$ models, whose distribution is depicted in Fig. 3 (additional results in Appendix E). These values confirm that set size differences are not derived from a small set of isolated outliers but from a considerably large group of samples that see their conformal set increase when using the ViT trained in a supervised manner. These results suggest that the strategies used to train foundation models yield better CP metrics, resulting in conformalized models that can be deployed more safely on critical scenarios. It is important to stress that this study is limited due to the different dataset scales used for training (i.e., ImageNet alone is insufficient to train a foundation model). Our goal, however, is to understand the conformalization properties of readily available pre-trained models, regardless of how they were pre-trained.
@@ 4.2.2 Impact under Distribution Shifts
The theoretical guarantees of the coverage for conformal prediction hold under the hypothesis that the calibration set and the test set are drawn from the same distribution, i.e., data exchangeability assumption. In this section, we analyze the impact of having calibration sets that present distributional shifts with regard to the testing set.
@@
We resort to ImageNet and its different versions: ImageNet-R [21], ImageNet-A [22], ImageNet-Sketch [62] and ImageNet-V2 [47]. To introduce the distributional drift between the calibration and testing data, we adapt the pre-trained model to one of the ImageNet versions. Then, ImageNet is used as the calibration set to conform the model, which is later tested on the ImageNet version used for adaptation. This is repeated for each ImageNet variant.
@@
APS [48] exhibits strong robustness against large distributional shifts, at the cost of substantially degrading efficiency. One would expect that adaptive CP methods, such as APS and RAPS, somehow mitigate domain shifts due to their adaptive nature. Nevertheless, Fig. 4 reveals several interesting observations, which contradict this intuition. First, we can observe that, when resorting to APS as CP method the coverage gap is consistently satisfied (or nearly satisfied) across all domain shifts and models (Fig. 4, middle). In contrast, RAPS generally shows very similar performances compared to LAC, obtaining lower marginal coverage under several models and domains, and substantially lower than APS. To understand this phenomenon, we now study how set sizes evolve across domains for the different methods (Fig. 4, left). We can easily observe that (i) APS yield the largest conformal sets across ImageNet domains, regardless of the model, and (ii) APS experiences the largest set increases when the complexity of the domain grows. Thus, as exposed in the previous Section, APS satisfied marginal coverage by substantially including more predicted classes, therefore increasing conformal set sizes.
@@
Figure 5: Domain shift analysis. Distribution of class-conditional coverages for CLIP on ImageNet-A: APS (left) and RAPS (right).
ECE ($\downarrow$) AvgSize ($\downarrow$) MCCC ($\uparrow$)
(×$\times 10^{-2}$) LAC APS RAPS LAC APS RAPS
@@
Table 2: Quantitative impact of calibrating vision foundation models. Average size and minimum class conditional coverage are reported between uncalibrated/calibrated models ($T=1$ and $T=1.1$ respectively) for several models, with ImageNet as benchmark dataset. Arrows indicate a decrease () or an increase() in metric after calibration, color indicates better (green) and worse (red) performance.
Figure 6: Impact of the temperature $T$ on ECE ($\downarrow$), set size ($\downarrow$), MCCC ($\uparrow$), and CovGap ($\downarrow$) (CLIP on ImageNet). Green region indicates the area with ECE smaller than for $T=1$. Green dotted line indicates value of optimal calibration ($T=1.1$). More plots in Appendix D.
APS and RAPS are adaptive methods that produce similar minimum class-conditional coverage, as exhibited in ImageNet-A (Fig. 4, right). However, they showcase significant differences in the average coverage gap. In the following, we explore this phenomenon in detail. More concretely, we depict in Fig. 5 the distribution of the conditional class coverage values obtained by APS and RAPS on ImageNet-A for CLIP (more datasets and models in Appendix F). Interestingly, while both approaches see their minimum class-conditional coverage decrease, their distributions are completely different. Indeed, APS distributions exhibit a Gaussian shape, with a decreasing number of categories presenting lower conditional coverage as they separate from 1-$\alpha$. In contrast, the distribution in RAPS exposes a significantly worse scenario, where class-conditional coverages are almost uniformly spread, with a non-negligible amount of classes below the expected coverage, i.e., $1-\alpha$.
4.2.3 Does calibration affect Conformal Prediction?
Confidence calibration is a popular strategy to improve the uncertainty estimates of deep models. These techniques, which can be either added as a post-processing step [44, 16, 29] or integrated as a training regularizer [19, 5, 38], adapt the model softmax predictions to reflect their actual performance accurately. The exponentially growing adoption of vision and vision-language foundation models in critical areas makes integrating confidence calibration a natural progression, as evidenced by recent works [36, 66, 59]. Thus, in this section, we investigate this important issue, as the relationship between calibration and conformal prediction in vision foundation models remains largely unexplored. Specifically, we examine whether calibrating these models with the popular Temperature Scaling (TS) [16] affects the conformal performance of fixed and adaptive CP methods. Particularly, we apply TS to the ImageNet results in the general case (Section 4.2.1). As a proper validation set is not available, we evaluate the CP performance over a set of $T$ values (14 values from $0.85$ to $2$), and found that $T=1.1$ typically yielded well-calibrated models [^1] Note that our goal is not to obtain the best-calibrated model but evaluate the impact of calibration on CP. Furthermore, this value aligns with existing works [34, 24] using TS for similar datasets, e.g., TinyImageNet..
Our observations suggest that calibration decreases the efficiency of CP sets but increases the minimum class-conditional coverage, improving the coverage gap.
Confidence calibration typically smooths the distribution of the class softmax scores, which results in less confident predictions. Consequently, the dominant value in the predicted softmax vector is typically lower in calibrated models. Nevertheless, since CP methods provide theoretical guarantees (under the data exchangeability assumption [61]) to satisfy the target marginal coverage of $1-\alpha$, these changes in the softmax distributions affect the conformalization obtained pre-TS. We present in Table 2 the results for the average set size and minimum class-conditional coverage before and after scaling the logits with TS. Furthermore, we include the Expected Calibration Error (ECE) values to verify that model calibration has improved. These results show that if the model is calibrated, its prediction set size tends to be larger, particularly for adaptive CP approaches (RAPS, and more specifically, APS). Fig. 7 further delves into these results, where we plot the distribution of differences between the set size of samples before and after calibration (i.e., a point in the distribution is $\mathcal{C}(x_{i})-\mathcal{C}(x_{i}^{TS})$). We can identify that the overall larger size in APS is caused by a consistent efficiency degradation across samples and not a few atypical cases with large conformal sets. We also observe that class-conditional coverage is typically improved on calibrated models, particularly when they are conformalized by adaptive CP methods. Thus, these results suggest that calibrating vision foundation models decreases efficiency while marginally enhancing class-conditional coverage, particularly on adaptive CP.
@@ 4.2.3 Does calibration affect Conformal Prediction?
The best coverage gap obtained by APS [48] approximately aligns with the coverage gap achieved at the optimal calibration point, whereas RAPS [1] coverage gap strongly differs. Fig. 6 shows the impact of the temperature $T$ on the different CP metrics on CLIP conformalized predictions. An interesting observation is that at the optimal temperature value $T=1.1$, the CovGap of APS is very close to its optimal point [^2] Results on Appendix D show similar behavior for other models. (0.0567 vs 0.0561). In contrast, while the RAPS coverage gap at optimal calibration is 0.0701, it decreases to 0.0642 as $T$ increases. This suggests that we could anticipate the best coverage gap for APS by assessing the model calibration performance, which will likely be near optimal calibration. On the downside, APS conformal sets efficiency is degraded. As $T$ increases, the softmax distributions get smoother, explaining the degradation in efficiency, which is even more drastic for APS, whose set size monotonically increases with $T$. We concede that, whereas APS minimizes the coverage gap, it does so at the expense of increasing the conformal set size. However, we believe that increasing the size, up to some extent, while improving coverage gap is preferable in critical decision systems.

### v3

@@ 4.2.1 Performance in the General Setting
First, we study the performance of 17 vision foundation models paired with CP methods under the standard setting, and on the three datasets, which present ideal conditions: a sufficiently large calibration set, and absence of distributional drifts between calibration and test sets. We aim to determine whether we can prescribe a winner solution in this scenario and which factors can help identify it.
@@
Figure 1 depicts the relationship between the linear probing performance of each model and the conformal set size (top), and coverage gap (bottom). The initial observation highlights a clear trend: higher-performance models tend to produce smaller prediction sets, regardless of the conformal method used. Nevertheless, while set efficiency (i.e. size) is typically considered as a sufficient condition in most prior literature in conformal prediction (Angelopoulos et al., 2020), our analysis revealed that the relation with the accuracy does not consistently hold when examining other metrics. In particular, Figure 1 (bottom) exposes that different CP methods yield mixed results for the coverage gap, which do not correlate directly with accuracy across all CP approaches. While APS appears to be almost unaffected by the model performance, RAPS clearly benefits from a strong performance, and LAC is negatively affected by more accurate models as the dataset becomes more complex.
@@
Regarding comparison between the CP methods, if we consider the set size, APS is clearly outperformed by the other approaches, whereas LAC provides the smallest prediction sizes, closely followed by RAPS. Indeed, RAPS is specifically designed to reduce the conformal set size of APS. However, the coverage gap results indicate that this comes at the cost of increasing the range of the class-conditional coverage. Below, we analyze the underlying causes that may explain this behaviour.
@@
RAPS class-conditional coverage, and therefore coverage gap, are more sensitive to the model’s accuracy. Let us assume we have two models, $M_{1}$ and $M_{2}$, in a multi-class classification problem, whose accuracies are $Acc_{1}$ and $Acc_{2}$, respectively. For each class $y\in\mathcal{Y}$, we refer to $C_{M_{i}}(y)=\mathbb{P}(Y\in\mathcal{S}_{M_{i}}(X)|Y=y)$ as the class-conditional coverage for $y$ under model $M_{i}$, which measures the probability that predictions include the true label when the true label is $y$. Furthermore, let $\delta_{M_{1}}=\min_{y\in\mathcal{Y}}C_{M_{1}}(y)$ and $\delta_{M_{2}}=\min_{y\in\mathcal{Y}}C_{M_{2}}(y)$ denote the minimum class-conditional coverage achieved by each model. Under this scenario, we argue that due to the penalty in RAPS, if $Acc_{1}<Acc_{2}$, then $\delta_{M_{1}}<\delta_{M_{2}}$ when using RAPS as a conformal prediction method. In particular, $M_{1}$ (with lower accuracy) needs to expand its prediction sets for certain classes to meet the marginal coverage target $1-\alpha$. However, the penalty term encourages small prediction sets, limiting an excessive number of classes. Thus, for some difficult classes, model $M_{1}$ may still potentially fail to meet the target coverage, as the enforced penalty discourages overly large sets. This ultimately results in lower coverage rates for those specific classes than $M_{2}$. In contrast, since APS does not include any regularization term that encourages small set sizes, it will compensate for lower performing models by increasing its set sizes, ultimately attaining higher class-conditional coverages.
@@
Figure 2: Comparison (APS vs RAPS) of the class-conditional coverage and set size for the class for which RAPS has the worst class-conditional coverage. Experiments performed on CIFAR-100. Models sorted (in ascending order) by their LP performance ($\min=0.64$ and $\max=0.92$), indicated by the size of the circles.
@@
To confirm this hypothesis, we perform the following experiment. First, we identify the class with the lowest minimum class-conditional coverage obtained by RAPS for each model $m$, which we refer to as $y^{R}_{m}$, and find its corresponding set size, both represented blue circles in Figure 2. Then, we identify the class-conditional coverage and set size of the class $y^{R}_{m}$ provided by APS across all models, whose values are shown as pink circles in both plots of Figure 2. Last, we also include the minimum class-conditional coverage obtained by APS across each method, depicted with a cross. Note that the minimum class-conditional coverage from APS does not necessarily correspond to the coverage of class $y^{R}_{m}$. Upon close examination of these results, we observe that, indeed, for models presenting lower accuracy (those with smaller circles), the gap between the class-conditional coverage for the worst class in RAPS and the same class in APS is consistently larger than for more accurate models (larger circles). This empirical evidence shows that the minimum class-conditional coverage is substantially reduced because RAPS is constrained from expanding its predictions set size. This effect is more pronounced in less accurate models, where the true class may rank far away from the maximum allowable set size in the softmax predictions. In contrast, APS tolerates worse models by increasing the set size, which ultimately degrades the set efficiency but yields better class-conditional coverages.
@@
Set size ($\downarrow$) MCCC ($\uparrow$) Coverage
Acc ($\uparrow$) LAC APS RAPS LAC APS RAPS LAC APS RAPS
$\text{ViT}_{\text{CLIP}}$ 72.01 3.03 9.50 3.73 0.434 0.556 0.418 0.900 0.900 0.900
$\text{ViT}_{\text{DINO-S}}$ 74.92 3.27 10.02 4.23 0.433 0.477 0.412 0.900 0.900 0.900
$\text{ViT}_{\text{MetaCLIP}}$ 75.80 2.39 9.43 2.84 0.479 0.535 0.467 0.900 0.900 0.900
$\text{ViT}_{\text{ImageNet}}$ 76.08 2.36 38.75 4.46 0.416 0.495 0.405 0.900 0.900 0.900
Following this analysis, we are also interested in determining whether a network pre-trained following a more traditional approach (i.e., standard supervised fine-tuning) offers similar conformal capabilities to self-supervised and contrastive ones. In particular, we select a ViT-B pre-trained on ImageNet, which is the same architecture as the visual encoder of the different foundation models. The results from this analysis (Table 1) reveal that, despite obtaining lower classification accuracy when using LP on the different foundation models, CP methods typically yield better performance than in $\text{ViT}_{\text{ImageNet}}$. This analysis should be put in perspective with the analysis in Figure 1, which found that higher accuracy leads to lower set size, indicating that the training scheme plays a very important role. These differences are significant under the APS approach, where the set size is significantly degraded on $\text{ViT}_{\text{ImageNet}}$. Moreover, the class-conditional coverage is also substantially affected, with nearly 6% decrease compared to the best model. Note that our criterion for model selection in Table 1 was to ensure comparable predictive capabilities, essential to enable a fair assessment of their conformal performance, as differences in accuracy could otherwise confound the analysis. As DINO-B classification performance is not comparable, conclusions from a conformalization standpoint cannot be drawn from it (see Table 7).
@@
To further delve into these differences, we compute, for each test sample, the difference between the conformal set size for APS when applied to $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{CLIP}}$ models, whose distribution is depicted in Figure 3 (additional results in Appendix Section D). These values confirm that set size differences are not derived from a small set of isolated outliers but from a considerably large group of samples that see their conformal set increase when using the ViT trained in a supervised manner. These results suggest that the strategies used to train foundation models yield better CP metrics, resulting in conformalized models that can be deployed more safely on critical scenarios. It is important to stress that this study is limited due to the different dataset scales used for training (i.e., ImageNet alone is insufficient to train a foundation model). Our goal, however, is to understand the conformalization properties of readily available pre-trained models, regardless of how they were pre-trained.
@@ 4.2.2 Impact under Distribution Shifts
The theoretical guarantees of the coverage for conformal prediction hold under the hypothesis that the calibration set and the test set are drawn from the same distribution, i.e., data exchangeability assumption. In this section, we analyze the impact of having calibration sets that present distributional shifts with respect to the testing set.
@@
We resort to ImageNet and its different versions: ImageNet-R (Hendrycks et al., 2021a), ImageNet-A (Hendrycks et al., 2021b), ImageNet-Sketch (Wang et al., 2019) and ImageNet-V2 (Recht et al., 2019). To introduce the distributional drift between the calibration and testing data, we adapt the pre-trained model to ImageNet. Then, ImageNet is used as the calibration set to conform the model, which is later tested on the ImageNet version used for adaptation. This is repeated for each ImageNet variant.
@@
APS (Romano et al., 2020) exhibits strong robustness against large distributional shifts, at the cost of substantially degrading efficiency. One would expect that adaptive CP methods, such as APS and RAPS, somehow mitigate domain shifts due to their adaptive nature. Nevertheless, Figure 5 reveals several interesting observations, which contradict this intuition. First, we can observe that, when resorting to APS as CP method the coverage gap is consistently satisfied (or nearly satisfied) across all domain shifts and models (Figure 5, middle). In contrast, RAPS generally shows very similar performances compared to LAC, obtaining lower marginal coverage under several models and domains, and substantially lower than APS. To understand this phenomenon, we now study how set sizes evolve across domains for the different methods (Figure 5, left). We can easily observe that (i) APS yield the largest conformal sets across ImageNet domains, regardless of the model, and (ii) APS experiences the largest set increases when the complexity of the domain grows.
@@
Figure 4: Domain shift analysis. Distribution of class-conditional coverages for CLIP (ViT-B) on ImageNet-A: APS (left) and RAPS (right).
Thus, as exposed in the previous Section, APS satisfied marginal coverage by substantially including more predicted classes, therefore increasing conformal set sizes. The regularization term in RAPS, which pushes towards lower set sizes, limits the adaptation to the domain shift, coming at the cost of a decrease in coverage.
Figure 5: Evaluation under domain-shift. Set size ($\downarrow$), coverage ($\uparrow$), and MCCC ($\uparrow$) across three CP methods and three foundation models. ImageNet versions are sorted based on OOD performance in Silva-Rodríguez et al. (2024).
APS and RAPS are adaptive methods that produce similar minimum class-conditional coverage, as exhibited in ImageNet-A (Figure 5, right). However, they showcase significant differences in the average coverage gap. In the following, we explore this phenomenon in detail. More concretely, we depict in Figure 4 the distribution of the conditional class coverage values obtained by APS and RAPS on ImageNet-A for CLIP (ViT-B), more datasets and models in Appendix Section E. Interestingly, while both approaches see their minimum class-conditional coverage decrease, their distributions are completely different. Indeed, APS distributions exhibit a Gaussian shape, with a decreasing number of categories presenting lower conditional coverage as they separate from 1-$\alpha$. In contrast, the distribution in RAPS exposes a significantly worse scenario, where class-conditional coverages are almost uniformly spread, with a non-negligible amount of classes below the expected coverage, i.e., $1-\alpha$.
4.2.3 Does calibration affect Conformal Prediction?
Confidence calibration is a popular strategy to improve the uncertainty estimates of deep models. These techniques, which can be either added as a post-processing step (Platt et al., 1999; Guo et al., 2017; Kull et al., 2019) or integrated as a training regularizer (Hebbalaguppe et al., 2022; Bohdal et al., 2023; Müller et al., 2020), adapt the model softmax predictions to reflect their actual performance accurately. The exponentially growing adoption of vision and vision-language foundation models in critical areas makes integrating confidence calibration a natural progression, as evidenced by recent works (Murugesan et al., 2024a; Yoon et al., 2024; Tu et al., 2024). Thus, in this section, we investigate this important issue, as the relationship between calibration and conformal prediction in vision foundation models remains largely unexplored. Specifically, we examine whether calibrating these models with the popular Temperature Scaling (TS) (Guo et al., 2017) affects the conformal performance of fixed and adaptive CP methods. Particularly, we apply TS to the ImageNet results in the general case (Section 4.2.1). As a proper validation set is not available, we evaluate the CP performance over a set of $T$ values (14 values from $0.85$ to $2$), and found that $T=1.1$ typically yielded well-calibrated models [^1] Note that our goal is not to obtain the best-calibrated model but evaluate the impact of calibration on CP. Furthermore, this value aligns with existing works (Mukhoti et al., 2020; Joy et al., 2023) using TS for similar datasets, e.g., TinyImageNet.. In Appendix Section C, we explore histogram binning as another method for model calibration.
Our observations suggest that calibration decreases the efficiency of CP sets, typically increasing the minimum class-conditional coverage, particularly on adaptive conformal prediction methods.
Effect on set efficiency. Confidence calibration typically smooths the distribution of the class softmax scores, which results in less confident predictions. Consequently, the dominant value in the predicted softmax vector is typically lower in calibrated models. Nevertheless, since CP methods provide theoretical guarantees (under the data exchangeability assumption (Vovk et al., 2005)) to satisfy the target marginal coverage of $1-\alpha$, these changes in the softmax distributions affect the conformalization obtained pre-TS. We present in Table 2 the results for the average set size and minimum class-conditional coverage before and after scaling the logits with TS. Furthermore, we include the Expected Calibration Error (ECE) values to verify that model calibration has improved.
These results show that if the model is calibrated, its prediction set size tends to be larger, particularly for adaptive CP approaches (RAPS, and more specifically, APS). Figure 6 further delves into these results, where we plot the distribution of differences between the set size of samples before and after calibration (i.e., a point in the distribution is $\mathcal{C}(x_{i})-\mathcal{C}(x_{i}^{TS})$). We can identify that the overall larger size in APS is caused by a consistent efficiency degradation across samples and not a few atypical cases with large conformal sets.
Effect on class-conditional coverage. We also observe that class-conditional coverage is typically improved on calibrated models, particularly when they are conformalized by adaptive CP methods. Thus, these results suggest that calibrating vision foundation models decreases efficiency while marginally enhancing class-conditional coverage, particularly on adaptive CP.
Table 2: Quantitative impact of calibrating vision foundation models with TS. Average size and minimum class-conditional coverage are reported between uncalibrated/calibrated models ($T=1$ and $T=1.1$ respectively) for several models, with ImageNet as benchmark dataset. Arrows indicate a decrease () or an increase() in metric after calibration, color indicates better (green) and worse (red) performance.
ECE ($\downarrow$) Set size ($\downarrow$) MCCC ($\uparrow$)
($\times 10^{-2}$) LAC APS RAPS LAC APS RAPS
@@ 4.2.3 Does calibration affect Conformal Prediction?
Effect on coverage gap. The best coverage gap obtained by APS (Romano et al., 2020) approximately aligns with the coverage gap achieved at the optimal calibration point, whereas RAPS (Angelopoulos et al., 2020) coverage gap strongly differs. Figure 7 shows the impact of the temperature $T$ on the different CP metrics on CLIP conformalized predictions. An interesting observation is that at the optimal temperature value $T=1.1$, the CovGap of APS is very close to its optimal point [^2] Results on Appendix Section C show similar behaviour for other models. (0.0567 vs 0.0561). In contrast, while the RAPS coverage gap at optimal calibration is 0.0701, it decreases to 0.0642 as $T$ increases. This suggests that the coverage gap for APS tends to be associated with model calibration performance, with smaller gaps occurring near optimal calibration. On the downside, APS conformal sets efficiency is degraded. As $T$ increases, the softmax distributions get smoother, explaining the degradation in efficiency, which is even more drastic for APS, whose set size monotonically increases with $T$. We concede that, whereas APS minimizes the coverage gap, it does so at the expense of increasing the conformal set size. However, we believe that increasing the size, up to some extent, while improving coverage gap is preferable in critical decision systems.
Figure 7: Impact of the temperature $T$ on ECE ($\downarrow$), set size ($\downarrow$), MCCC ($\uparrow$), and CovGap ($\downarrow$) (CLIP (ViT-B) on ImageNet). Green region indicates the area with ECE smaller than for $T=1$. Green dotted line indicates value of optimal calibration ($T=1.1$). More plots in Appendix Section C.